# Quantum information with quantum-like bits

### v5

@@ I Introduction
The present noisy-intermediate scale quantum (NISQ) era of quantum devices is witnessing the successful implementation and test of diverse quantum algorithms on medium-size quantum hardware, from search on unsorted databases, Zhang et al. (2021) to variational quantum algorithms, Cerezo et al. (2021) to simulations of open quantum dynamics. Sun, Shih, and Cheng (2024) However, hardware of at least hundreds of thousands of physical qubits, way beyond the current engineering capabilities, is needed to obtain practical quantum advantage. This involves storing and processing the information volume required by the most challenging and relevant computational problems. Beverland et al. (2022) The critical issue limiting the scalability of quantum resources is the buildup of uncontrolled noise and dissipation, leading to loss of coherence and fidelity. Preskill (2018); Lau et al. (2022); Gill et al. (2022)
@@
A large class of regular graphs encodes the correlations required to identify robust collective states. This general feature can be exploited for a variety of applications. For example, we discussed in Ref. Scholes, 2025a how this principle can support the engineering of molecular delocalization. The same reference, along with Refs. Krebs and Shaheen, 2011; Brouwer and Haemers, 2011, offers the interested reader background information on the mathematical formalism of regular and expander graphs, as well as their spectral properties. In the spirit of quantum-inspired computing, we showed in recent work that network correlations can serve as resources for quantum information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2025) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??. Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red). A spectral analysis of this class of graphs highlights the presence of two isolated emergent states. These an can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??). The robustness of emergent states against noise appears to be a direct consequence of synchronization. Scholes (2023, 2020) This intriguing feature is related to the occurrence of classical synchronization phenomena. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref. Scholes, 2024 that the emergent states of connected expander graphs are isomorphic to the Hilbert space of a qubit. This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks. In Ref. Scholes and Amati, 2025 we built on these results and proposed a general framework to construct resources for multiple QL bits. In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
@@
Our approach leads to a natural correspondence between “QL gates” transformations and the steady states of the classical oscillatory dynamics. Additionally, we propose a variety of experimental resources that can be utilized to test the formalism. This supports the applicability of the approach for information protocols involving low-dimensional Hilbert spaces, where the exponentially large complexity of the classical phase space can be handled by available experimental resources. Finally, we connect our perspective to the QL probabilistic framework developed by Khrennikov. Khrennikov (2007a, 2019, 2010) This theory formalizes the notion that genuine quantum systems are not required to process information in a QL way. Instead, to observe nonclassical interference it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and projective measurements.
The present paper is structured as follows. In ?? we summarize our previous work describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we show the fundamental role of emergent states in the synchronizing dynamics of the Kuramoto model as a paradigmatic classical system of synchronizing oscillators. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to implement and verify experimentally our formalism.
@@ II Quantum-like bits
In this section we derive a mapping from the emergent states of classical networks to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. The present analysis provides a complete description of the axioms of states and operators in quantum mechanics, within the notation of the emergent states. Later in the paper we show the direct connection between this abstract approach and the dynamics of many-body classical systems.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents. The pair-wise interactions can be mapped onto two graphs $G_{1}$ and $G_{2}$. Here, we assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. Also, we consider at this stage undirected simple graphs, without loops (while in the paper we will discuss graph transformations that break these symmetries). The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
@@
and $n=1,2$. A key property of these subgraphs is that they exhibit an emergent state—a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) The direct sum of the two adjacency matrices is
@@
involving a coupling matrix $C$. Note that $C$ does not represent a simple graph by itself, i.e. it is in general non-symmetric, and it can admit diagonal entries connecting corresponding vertices in $A_{1}$ and $A_{2}$. Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph. We sample $C$ by randomly assigning $1$ to $N_{\mathrm{C}}\leq{N_{G}}^{2}$ of its entries. This matrix is chosen to be regular with valency $l=N_{\mathrm{C}}/{N_{G}}$ [similarly to ??]. ?? represents the fundamental resource used to build an isolated QL bit. Scholes (2024)
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. By mapping the states associated to those peaks onto a Hilbert space, we obtain a platform for information processing that can be controlled by network manipulation. We will elaborate on these important concepts further in the manuscript.
We study the spectral properties of the resources by plotting in ?? the distribution $\rho(\lambda)$ of the eigenvalues of ?? for increasing values of the valency $l$ of $C$.
Figure 3: Main panel: distribution of the eigenvalues $\lambda$ of the QL-resource ??, for increasing values of the order $l$ of the coupling matrix $C$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). Inset: largest and second largest eigenvalues of the spectrum as a function of $l$, with the same color code as in the main panel. We fix here ${N_{G}}=100$ and $k=25$.
Results are shown from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). The other parameters are fixed to ${N_{G}}=100$ and $k=25$. The emergent states can be identified for each $l$ from the isolated sharp peaks in the spectrum. For $l=0$ we observe a single isolated line associated to the two-fold degenerate eigenvalue $\lambda=k=25$ (red markers). As the valency $l$ increases, this emergent eigenvalue split into two, and the peaks shift to $\lambda_{\pm}=k\pm l$. We track the “ballistic” evolutions of the eigenvectors as a function of the coupling valency $l$ in the inset of ??. Here, the largest and second largest eigenvalues are shown as squared and dotted markers, respectively. While the value of the largest eigenvalue increases linearly for arbitrary $l$, we observe that the second largest eigenvalue plateaus at values close to $\tilde{\lambda}=2\sqrt{k-1}$. This barrier is defined by the radius of the Wigner semicircle distribution for a graph of order $k$, shown as a black dashed line in the figure. McKay (1981); Farkas et al. (2001) This portion of the spectrum is populated by random low-frequency states, whose high density makes them hardly distinguishable. Scholes (2020, 2023, 2024) From the present analysis we infer that, by fixing $l=(k-\tilde{\lambda})/2$, both the distance between the two emergent eigenvalues and between those and the dark states are maximized. As elucidated in Ref. Scholes, 2023, 2024, this analysis offers valuable insight for creating well-resolved and easily distinguishable emergent states for quantum information purposes.
@@
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights. We keep, however, the notation general, to account for possible structural disorder in the network leading to this break of symmetry. The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ?? with eigenvalues $\{k-l,k+l\}$, respectively. Note that in prior work we used the intention of the QL bit construction to represent explicitly the basis of the the subgraphs, so that the emergent state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and similarly $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ?? describes another basis choice, equally valid. The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??. In the following we show that the mapping in ?? leads to a well-defined Hilbert space. First of all, it is straightforward to show that the $\psi_{\sigma}$’s are mutually orthogonal and normalized:
$\braket{\psi_{\sigma},\psi_{\sigma^{\prime}}}$ $=\sum_{i=1}^{2{N_{G}}}\psi_{\sigma,i}^{*}\psi_{\sigma^{\prime},i}=\frac{1}{2}\left(|a_{1}|^{2}+(2\delta_{\sigma\sigma^{\prime}}-1)|a_{2}|^{2}\right)$
$=\delta_{\sigma\sigma^{\prime}}\mapsto\braket{{\sigma}{\sigma^{\prime}}}{,}$ (II.5)
where $\sigma,\sigma^{\prime}\in\mathcal{S}$. Operators acting on ?? are expressed in terms of the matrices in $M_{2{N_{G}}}(\mathbb{R})$
$\Psi_{\sigma\sigma^{\prime}}=\psi_{\sigma}\otimes\psi_{\sigma^{\prime}}^{\dagger}\mapsto\ket{\sigma}\bra{\sigma^{\prime}}.$ | . (II.6)
As it would be expected in an abstract Hilbert space, the maps in ?? combine according to the rule
$\Psi_{\sigma\sigma^{\prime}}\Psi_{\pi\pi^{\prime}}=\Psi_{\sigma\pi^{\prime}}\braket{\psi_{\sigma^{\prime}},\psi_{\pi}},$ (II.7)
and they transform under Hermitian conjugation as $\Psi_{\sigma\sigma^{\prime}}^{\dagger}=\Psi_{\sigma^{\prime}\sigma}$.
@@ II.2.1 Formalism
Various representations of the quantum states are possible with respect to classical network structures. For a small number of QL bits the representation defined in Ref. Scholes, 2024 is useful. In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of entanglement. In Ref. Scholes and Amati, 2025 we propose a formal representation of states applicable to an arbitrary number of QL bits. This is based on the Cartesian product of single-QL-bit graphs. Consistently with that work, we define here as a simple generalization of ?? the resource for a system of ${N_{\mathrm{QL}}}$ bits
$\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.8)
@@
$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}},$ (II.10)
are eigenvectors of ??. Here, different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\cdots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$. The components of ?? are expressed via the multi-index $\bm{i}=\left\{i^{(1)},\cdots,i^{(N_{\mathrm{QL}})}\right\}$ as
$[\psi_{\bm{\sigma}}]_{\bm{i}}$ $=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}.$ (II.11)
As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\mathcal{S}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@
$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.12)
@@
$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.13)
@@
In analogy to ?? for ${N_{\mathrm{QL}}}=1$, the states in ?? define an orthonormal basis for $\mathcal{H}$ according to the scalar product defined by
@@
$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.14)
Similarly, we extend the mapping of operators in ?? to
$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\otimes\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.15)
with components
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}\right]_{\bm{i,j}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},i^{(q)}}\psi^{*}_{\sigma^{\prime(q)},j^{(q)}}.$ (II.16)
The Hermitian conjugate of ?? is defined such that
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.17)
and we generalize ?? to
$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.18)
The mapping discussed here provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits. In the next ?? we show explicitly, as an example, how to construct maximally entangled Bell states for ${N_{\mathrm{QL}}}=2$.

### v6

@@ I Introduction
The present noisy-intermediate scale quantum era of quantum devices is witnessing the successful implementation and test of diverse quantum algorithms on medium-size quantum hardware, from search on unsorted databases, Zhang et al. (2021) to variational quantum algorithms, Cerezo et al. (2021) to simulations of open quantum dynamics. Sun, Shih, and Cheng (2024) However, hardware of at least hundreds of thousands of physical qubits, way beyond the current engineering capabilities, is needed to obtain practical quantum advantage. This involves storing and processing the information volume required by the most challenging and relevant computational problems. Beverland et al. (2022) The critical issue limiting the scalability of quantum resources is the buildup of uncontrolled noise and dissipation, leading to loss of coherence and fidelity. Preskill (2018); Lau et al. (2022); Gill et al. (2022)
@@
A large class of regular graphs encodes the correlations required to identify robust collective states. This general feature can be exploited for a variety of applications. For example, we discussed in Ref. Scholes, 2025 how this principle can support the engineering of molecular delocalization. The same reference, along with Refs. Krebs and Shaheen, 2011; Brouwer and Haemers, 2011, offers the interested reader background information on the mathematical formalism of regular and expander graphs, as well as their spectral properties. In the spirit of quantum-inspired computing, we showed in recent work that network correlations can serve as resources for quantum information processing. Scholes (2020, 2023, 2024); Scholes and Amati (2025) We focused, in particular, on many body systems whose correlations obey the graph structure shown in panel a) of ??. Here, two densely connected regions (in blue) of a network are mutually interacting via sparse links (in red). A spectral analysis of this class of graphs highlights the presence of two isolated emergent states. These can be clearly distinguished from a broad band of states arising from random network configurations (dashed lines in the spectra of ??). The robustness of emergent states against noise appears to be a direct consequence of synchronization. Scholes (2023, 2020) This intriguing feature is related to the occurrence of classical synchronization phenomena. Haken (1975); Fano (1992); Strogatz (2012); Artime and De Domenico (2022) We showed in Ref. Scholes, 2024 that the emergent states of connected expander graphs are isomorphic to the Hilbert space of a qubit. This led us to introducing the notion of quantum-like (QL) bits within the context of classically-synchronizing networks. In Ref. Scholes and Amati, 2025 we built on these results and proposed a general framework to construct resources for multiple QL bits. In particular, we showed that the Cartesian product ($\square$) of ${N_{\mathrm{QL}}}$ bipartite graphs obeys the same dimensionality of a register of ${N_{\mathrm{QL}}}$ qubits [see panel b) of ??].
@@
Our approach leads to a natural correspondence between “QL gates” transformations and the steady states of the classical oscillatory dynamics. Additionally, we propose a variety of experimental resources that can be utilized to test the formalism. This supports the applicability of the approach for information protocols involving low-dimensional Hilbert spaces, where the exponentially large complexity of the classical phase space can be handled by available experimental resources. Finally, we connect our perspective to the QL probabilistic framework developed by Khrennikov. Khrennikov (2007a, 2019, 2010) This theory formalizes the notion that genuine quantum systems are not required to process information in a QL way. Instead, to observe nonclassical interference it is sufficient to construct a contextual probability theory involving suitable notions of QL states, observables, and projective measurements. This paper can be read alongside other recent developments by the authors within the same framework, as presented in the complementary work in Ref. Amati and Scholes, 2025. In that study, we show that arbitrary quantum states can be encoded into the synchronizing steady state of a nonlinear oscillator system. We also apply the approach proposed here to construct a QL analog of entanglement. Although the two works are closely related, each is self-contained and can be read independently.
The present paper is structured as follows. In ?? we summarize our previous work describing how to construct a single-QL-bit Hilbert space. In ?? we discuss its generalization to higher dimensions, and we show explicitly in ?? how to construct Bell states for two QL bits. In ?? we show the fundamental role of emergent states in the synchronizing dynamics of the Kuramoto model as a paradigmatic classical system of synchronizing oscillators. In ?? we study the action of a complete set of QL gates, and we provide concrete examples of their action. In ??, we present a set of axioms that enables us to define a QL probabilistic theory in our approach. In particular, we highlight the presence of a nonclassical interference term in the relation between total and conditional probabilities, specific to non-Kolmogorovian contextual probabilistic models. Khrennikov (2010) Finally, in ??, we propose a variety of experimental resources that could be utilized to experimentally implement and verify our formalism.
@@ II Quantum-like bits
In this section we derive a mapping from the emergent states of classical networks to the Hilbert space of interacting QL bits. While the concept has been already elucidated in Ref. Scholes, 2024, here we lay out a more formal description. The present analysis provides a complete description of the axioms of states and operators in quantum mechanics, within the notation of the emergent states. Later in the paper we show the direct connection between this abstract approach and the dynamics of classical nonlinear oscillators.
@@ II.1 Isolated system
Let us consider two classical many-body systems with pairwise interactions between the elementary constituents mapped onto two graphs, $G_{1}$ and $G_{2}$. Here, we assume here that both graphs involve the same number of vertices, ${N_{G}}$, and we label by $A_{1}$ and $A_{2}$ the related adjacency matrices. Also, we consider at this stage undirected simple graphs, without loops (while in the paper we will discuss graph transformations that break these symmetries). The adjacency matrix for this class of graphs is symmetric with entries either 00 or $1$ and null diagonal. Additionally, we assume an isotropic structure for all two-body interactions, such that the valence of the graphs is constant and equal to $k$. Diestel (2017) In particular,
@@
and $n=1,2$. The direct sum of the two adjacency matrices is
@@
involving a coupling matrix $C$. Note that $C$ does not represent a simple graph by itself, i.e. it is in general non-symmetric, and it can admit diagonal entries connecting corresponding vertices in $A_{1}$ and $A_{2}$. Conversely, ??, being overall symmetric and loop-free, describes a simple undirected graph. The matrix $C$ is also chosen to be regular, with valency $l$. ?? represents the fundamental resource used to build an isolated QL bit. Scholes (2024)
@@
As discussed in Refs. Scholes, 2023, 2024, resources with structure as in ?? are particularly appealing for information processing. In fact, these resources allow for emergent states associated to distinguishable eigenvalues, well isolated from a large band of random states. The occurrence of emergent states is a common feature of Erdös-Renyi random graphs and expander families, such as the k-regular graphs. Scholes, DelPo, and Kudisch (2020); Scholes (2024); Lubotzky (2012) By mapping the emergent states onto a Hilbert space, we obtain a platform for information processing that can be controlled by suitable network processing. We will elaborate on these important concepts further in the manuscript.
We study the spectral properties of the graph resources by plotting in ?? the distribution $\rho(\lambda)$ of the eigenvalues of ?? for increasing values of the valency $l$ of the regular off-diagonal block $C$.
Figure 3: Main panel: distribution of the eigenvalues $\lambda$ of the QL resource ??, for increasing values of the order $l$ of the coupling matrix $C$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). Inset: largest and second largest eigenvalues of the spectrum as a function of $l$, with the same color code as in the main panel. We fix here ${N_{G}}=100$ and $k=25$.
More specifically, we sample a set of graphs with the structure given in ??. Given that each block of $\mathcal{R}$ defines a regular random graph, the only constraints given in the sampling are that each of the vertices in the $A_{i}$’s and $C$ is connected to $k$ and $l$ other vertices, respectively. In ?? we show a family of distributions for a fixed value of $k=25$ and increasing vaues of $l$, from $l=0$ (blue/dark) to $l={N_{G}}$ (yellow/light). Here, the sizes of all blocks are fixed to ${N_{G}}=100$. The emergent states can be identified for each $l$ from the couples of isolated sharp peaks (with identical colors), displaced symmetrically around $\lambda=k=25$. More specifically, for $l=0$ we observe a single isolated line associated to the two-fold degenerate eigenvalue $\lambda=k=25$ (red markers). As the valency $l$ increases, this emergent eigenvalue split into two, and the peaks shift to $\lambda_{\pm}=k\pm l$. We track the “ballistic” evolutions of the eigenvectors as a function of the coupling valency $l$ in the inset of ??. Here, the largest and second largest eigenvalues are shown as squared and dotted markers, respectively. While the value of the largest eigenvalue increases linearly for arbitrary $l$, we observe that the second largest eigenvalue plateaus at values close to $\tilde{\lambda}=2\sqrt{k-1}$. This barrier is defined by the radius of the Wigner semicircle distribution for a graph of order $k$, shown as a black dashed line in the figure. McKay (1981); Farkas et al. (2001) This portion of the spectrum is populated by random low-frequency states, whose high density makes them hardly distinguishable. Scholes (2020, 2023, 2024) From the present analysis we infer that, by fixing $l=(k-\tilde{\lambda})/2$, both the distance between the two emergent eigenvalues and between those and the dark states are maximized. As elucidated in Ref. Scholes, 2023, 2024, this analysis offers valuable insight for creating well-resolved and easily distinguishable emergent states for quantum information purposes.
@@
where $a_{i}$ is the normalized eigenvector associated to the largest eigenvalue of $A_{i}$. As both the $A_{i}$’s and $C$ are exactly regular, then $a_{i}=(1,1\ldots,1)^{\top}/\sqrt{2{N_{G}}}$ are vectors with constant weights. Here and in the following we decided, however, to keep a more general notation, to account for possible structural disorder in the network that could break this symmetry. The vectors $\psi_{\sigma}$, $\sigma\in\{\downarrow,\uparrow\}\equiv\mathcal{S}$, are eigenvectors of ?? with eigenvalues $\{k-l,k+l\}$, respectively. In earlier work, we used the QL bit construction to explicitly associate the subgraph basis with the emergent states: for example, the state $a_{1}$ of the graph with adjacency matrix $A_{1}$ represents $\ket{\downarrow}$, and $a_{2}$ represents $\ket{\uparrow}$.Scholes (2024); Scholes and Amati (2025) ?? presents a different, but equally valid, choice of basis.
The present approach establishes a direct connection between the computational basis and the emergent states of the composite graph described by ??.
@@ II.2.1 Formalism
Various representations of quantum states are possible with respect to classical network structures. For a small number of QL bits the representation defined in Ref. Scholes, 2024 is useful. In that work the Bell states, involving two QL bits, are designed to give insight on proof-of-principle examples of QL entanglement. In Ref. Scholes and Amati, 2025 we propose a formal representation of states applicable to an arbitrary number of QL bits. This is based on the Cartesian product of single-QL-bit graphs. Consistently with that work, we define here as a simple generalization of ?? the resource for a system of ${N_{\mathrm{QL}}}$ bits
$\mathcal{R}=\sum_{q=1}^{N_{\mathrm{QL}}}\left[\bigotimes_{p=1}^{q-1}\mathbb{1}_{2{N_{G}}}\right]\otimes\mathcal{R}^{(q)}\otimes\left[\bigotimes_{p^{\prime}=q+1}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}\right],$ (II.5)
@@
$\psi_{\bm{\sigma}}=\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\psi_{\sigma^{(q)}}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{\sigma^{(q)}}=\ket{\bm{\sigma}}$ (II.7)
are eigenvectors of ??. Different basis elements are labeled using the multi-index $\bm{\sigma}=\left\{\sigma^{(1)},\ldots,\sigma^{(N_{\mathrm{QL}})}\right\}\in\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}=\mathcal{S}^{N_{\mathrm{QL}}}$. As discussed after ??, each $\psi_{\sigma^{(q)}}$, with $\sigma^{(q)}\in\mathcal{S}$, is associated to one of the eigenvalues $k\pm l$ of $\mathcal{R}^{(q)}$. Let us label this eigenvalue $\lambda_{\sigma^{(q)}}$. The action of $\mathcal{R}$ on $\psi_{\bm{\sigma}}$ is then given by
@@
$=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}\psi_{\bm{\sigma}}=\lambda_{\bm{\sigma}}\psi_{\bm{\sigma}},$ (II.8)
@@
$\lambda_{\bm{\sigma}}=\sum_{q=1}^{N_{\mathrm{QL}}}\lambda_{\sigma^{(q)}}^{(q)}.$ (II.9)
@@
The states in ?? define an orthonormal basis for $\mathcal{H}$ according to the scalar product defined by
@@
$=\prod_{q=1}^{{N_{\mathrm{QL}}}}\delta_{{\sigma}^{(q)},{\sigma^{\prime}}^{(q)}}=\prod_{q=1}^{{N_{\mathrm{QL}}}}\braket{\sigma^{(q)}}{{\sigma^{\prime}}^{(q)}}.$ (II.10)
Elementary operators in our approach are defined by
$\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}$ $=\psi_{\bm{\sigma}}\psi_{\bm{\sigma}^{\prime}}^{\dagger}\mapsto\bigotimes_{q=1}^{{N_{\mathrm{QL}}}}\ket{{\sigma}_{q}^{(q)}}\bra{{\sigma^{\prime}}_{q}^{(q)}}=\ket{\bm{\sigma}}\bra{\bm{\sigma}^{\prime}},$ (II.11)
The Hermitian conjugates of ?? are defined such that
$\left[\Psi_{\bm{\sigma},\bm{\sigma}^{\prime}}^{\dagger}\right]_{\bm{ij}}=\prod_{q=1}^{N_{\mathrm{QL}}}\psi_{\sigma^{(q)},j^{(q)}}^{*}\psi_{\sigma^{\prime(q)},i^{(q)}}=\left[\Psi_{\bm{\sigma}^{\prime},\bm{\sigma}}\right]_{\bm{ij}},$ (II.12)
and they obey the orthogonality relations
$\Psi_{\bm{\sigma\sigma^{\prime}}}\Psi_{\bm{\pi\pi^{\prime}}}=\Psi_{\bm{\sigma\pi^{\prime}}}\braket{\bm{\psi_{\sigma^{\prime}}},\bm{\psi_{\pi}}}.$ (II.13)
Let us note that the Cartesian product of ${N_{\mathrm{QL}}}$ QL-bits, each defined by a graph with $2{N_{G}}$ vertices, yields a graph whose adjacency matrix has dimension $(2{N_{G}})^{N_{\mathrm{QL}}}$, scaling exponentially with the number of QL-bits. This behavior is, to some extent, an expected trade-off when encoding quantum states into classical many-body systems. As discussed in ??, this issue presents a practical—though not fundamental—limitation to testing the formalism with available computational and experimental resources.
The mapping proposed in this section provides a general description of state vectors, entanglement and operators constructed from the emergent states of an arbitrary number of QL bits. In the next ?? we show explicitly, as an example, how to construct maximally entangled Bell states for ${N_{\mathrm{QL}}}=2$.